OpenStudy (anonymous):

A jet flying due north at 8km/min passes 2km directly above a plane flying due wast at 4km/min. How quickly are they separating when the jet is 16km away from the crossover.

OpenStudy (anonymous):

Did I draw this correctly? 

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks good to me

OpenStudy (misty1212):

then pythagoras finishes it

OpenStudy (misty1212):

well pythagoras and a bunch of other calculations ...

OpenStudy (anonymous):

I'm not getting the correct answer :(

OpenStudy (freckles):

can i see your equations you used? And what you got after differentiating? And your final solution.

OpenStudy (anonymous):

I'm not sure if I did it right, but here it goes... I first find x and the answer was 6sqrt7 m by using x^2+y^2=D^2 then I use this equation: 2x(dx/dt)+ 2y(dy/dt)=2D(dD/dt) and sub in all the values in 2(6sqrt7)(8)+0=2(16)(dD/dt) and final answer is 96sqrt7/32, which is not the correct answer. (Correct answer is 80/9km/min)

OpenStudy (freckles):

why is 2y(dy/dt)=0?

OpenStudy (freckles):

in your picture you said dy/dt=4

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

I was thinking of a different scenerio on another question

OpenStudy (freckles):

\[2(6 \sqrt{7})8+2(2)(4)=2(16) \frac{dD}{dt}\]

OpenStudy (freckles):